OpenStudy (anonymous):

Which of the following systems of equations has no solution? Answer 9x + 5y = 1; 15y = 18x − 4 -7x − 7 = 3y; -14y − 8 = -6x 4x − 3y = 9; 6y = 8x − 18 7y = 5x − 10; 10x − 14y = 8

OpenStudy (cwrw238):

see if any of equation in a system is a multiple of the other if so that your answer

OpenStudy (cwrw238):

you'll need to rearrange in order to make is obvious

OpenStudy (cwrw238):

for example 2nd option write in form ax + by = n -7x - 3y = 7 -6x + 14y = -8 one is not multiply by other

OpenStudy (cwrw238):

do you follow that?

OpenStudy (anonymous):

Another reasonably quick way is to calculate the determinant. If it's not equal to 0, there is a solution.

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

and determinant of a equation  and if